\begin{thebibliography}{16}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Artetxe et~al.(2018)Artetxe, Labaka, and Agirre]{artetxe2018acl}
Mikel Artetxe, Gorka Labaka, and Eneko Agirre.
\newblock A robust self-learning method for fully unsupervised cross-lingual
  mappings of word embeddings.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pp.\  789--798, 2018.

\bibitem[Bojanowski et~al.(2016)Bojanowski, Grave, Joulin, and
  Mikolov]{DBLP:journals/corr/BojanowskiGJM16}
Piotr Bojanowski, Edouard Grave, Armand Joulin, and Tomas Mikolov.
\newblock Enriching word vectors with subword information.
\newblock \emph{CoRR}, abs/1607.04606, 2016.
\newblock URL \url{http://arxiv.org/abs/1607.04606}.

\bibitem[Cho et~al.(2014)Cho, van Merrienboer, Gulcehre, Bahdanau, Bougares,
  Schwenk, and Bengio]{Cho_2014}
Kyunghyun Cho, Bart van Merrienboer, Caglar Gulcehre, Dzmitry Bahdanau, Fethi
  Bougares, Holger Schwenk, and Yoshua Bengio.
\newblock Learning phrase representations using rnn encoder–decoder for
  statistical machine translation.
\newblock \emph{Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, 2014.
\newblock \doi{10.3115/v1/d14-1179}.
\newblock URL \url{http://dx.doi.org/10.3115/v1/D14-1179}.

\bibitem[Christodouloupoulos \& Steedman(2015)Christodouloupoulos and
  Steedman]{christodouloupoulos2015massively}
Christos Christodouloupoulos and Mark Steedman.
\newblock A massively parallel corpus: the bible in 100 languages.
\newblock \emph{Language resources and evaluation}, 49\penalty0 (2):\penalty0
  375--395, 2015.

\bibitem[Coates \& Bollegala(2018)Coates and
  Bollegala]{coates-bollegala-2018-frustratingly}
Joshua Coates and Danushka Bollegala.
\newblock Frustratingly easy meta-embedding {--} computing meta-embeddings by
  averaging source word embeddings.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, pp.\  194--198, New Orleans,
  Louisiana, June 2018. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/N18-2031}.
\newblock URL \url{https://www.aclweb.org/anthology/N18-2031}.

\bibitem[Grave et~al.(2018)Grave, Bojanowski, Gupta, Joulin, and
  Mikolov]{grave2018learning}
Edouard Grave, Piotr Bojanowski, Prakhar Gupta, Armand Joulin, and Tomas
  Mikolov.
\newblock Learning word vectors for 157 languages.
\newblock In \emph{Proceedings of the International Conference on Language
  Resources and Evaluation (LREC 2018)}, 2018.

\bibitem[Hill et~al.(2016)Hill, Cho, Korhonen, and Bengio]{TACL711}
Felix Hill, KyungHyun Cho, Anna Korhonen, and Yoshua Bengio.
\newblock Learning to understand phrases by embedding the dictionary.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  4\penalty0 (0):\penalty0 17--30, 2016.
\newblock ISSN 2307-387X.
\newblock URL \url{https://transacl.org/ojs/index.php/tacl/article/view/711}.

\bibitem[{isiXhosa Oxford Living Dictionary}(2020)]{oxford}
{isiXhosa Oxford Living Dictionary}.
\newblock {isiXhosa Oxford Living Dictionary}, 2020.
\newblock URL \url{https://xh.oxforddictionaries.com/}.

\bibitem[Klein et~al.(2017)Klein, Kim, Deng, Senellart, and
  Rush]{klein-etal-2017-opennmt}
Guillaume Klein, Yoon Kim, Yuntian Deng, Jean Senellart, and Alexander Rush.
\newblock {O}pen{NMT}: Open-source toolkit for neural machine translation.
\newblock In \emph{Proceedings of {ACL} 2017, System Demonstrations}, pp.\
  67--72, Vancouver, Canada, July 2017. Association for Computational
  Linguistics.
\newblock URL \url{https://www.aclweb.org/anthology/P17-4012}.

\bibitem[Lazaridou et~al.(2017)Lazaridou, Marelli, and
  Baroni]{doi:10.1111/cogs.12481}
Angeliki Lazaridou, Marco Marelli, and Marco Baroni.
\newblock Multimodal word meaning induction from minimal exposure to natural
  text.
\newblock \emph{Cognitive Science}, 41\penalty0 (S4):\penalty0 677--705, 2017.
\newblock \doi{10.1111/cogs.12481}.
\newblock URL \url{https://onlinelibrary.wiley.com/doi/abs/10.1111/cogs.12481}.

\bibitem[Mikolov et~al.(2013)Mikolov, Sutskever, Chen, Corrado, and
  Dean]{mikolov2013distributed}
Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg~S Corrado, and Jeff Dean.
\newblock Distributed representations of words and phrases and their
  compositionality.
\newblock In \emph{Advances in neural information processing systems}, pp.\
  3111--3119, 2013.

\bibitem[Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu.
\newblock Bleu: a method for automatic evaluation of machine translation.
\newblock In \emph{Proceedings of the 40th annual meeting on association for
  computational linguistics}, pp.\  311--318. Association for Computational
  Linguistics, 2002.

\bibitem[Pennington et~al.(2014)Pennington, Socher, and
  Manning]{pennington-etal-2014-glove}
Jeffrey Pennington, Richard Socher, and Christopher Manning.
\newblock {G}love: Global vectors for word representation.
\newblock In \emph{Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing ({EMNLP})}, pp.\  1532--1543, Doha, Qatar,
  October 2014. Association for Computational Linguistics.
\newblock \doi{10.3115/v1/D14-1162}.
\newblock URL \url{https://www.aclweb.org/anthology/D14-1162}.

\bibitem[Ruder(2017)]{DBLP:journals/corr/Ruder17}
Sebastian Ruder.
\newblock A survey of cross-lingual embedding models.
\newblock \emph{CoRR}, abs/1706.04902, 2017.
\newblock URL \url{http://arxiv.org/abs/1706.04902}.

\bibitem[Shikali et~al.(2019)Shikali, Sijie, Qihe, and
  Mokhosi]{shikali2019better}
Casper~S Shikali, Zhou Sijie, Liu Qihe, and Refuoe Mokhosi.
\newblock Better word representation vectors using syllabic alphabet: A case
  study of swahili.
\newblock \emph{Applied Sciences}, 9\penalty0 (18):\penalty0 3648, 2019.

\bibitem[Stanojevi{\'c} \& Sima{'}an(2014)Stanojevi{\'c} and
  Sima{'}an]{stanojevic-simaan-2014-beer}
Milo{\v{s}} Stanojevi{\'c} and Khalil Sima{'}an.
\newblock {BEER}: {BE}tter evaluation as ranking.
\newblock In \emph{Proceedings of the Ninth Workshop on Statistical Machine
  Translation}, pp.\  414--419, Baltimore, Maryland, USA, June 2014.
  Association for Computational Linguistics.
\newblock \doi{10.3115/v1/W14-3354}.
\newblock URL \url{https://www.aclweb.org/anthology/W14-3354}.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2020_conference,times}
\usepackage{tabularx}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\def\sD{{\mathbb{D}}}

\def\sV{{\mathbb{V}}}

\usepackage{booktabs}
\usepackage{tipa}

\usepackage{hyperref}
\usepackage{url}

\title{{Combining Pretrained High-Resource Embeddings and Subword Representations for Low-Resource Languages}}

\iclrfinalcopy
\author{Machel Reid, Edison Marrese-Taylor \& Yutaka Matsuo\\
Graduate School of Engineering\\
The University of Tokyo\\
Tokyo, Japan \\
\texttt{machelreid2004@gmail.com, \{emarrese,matsuo\}@weblab.t.u-tokyo.ac.jp} \\
}

\begin{document}

\maketitle

\begin{abstract}
The contrast between the need for large amounts of data for current Natural Language Processing (NLP) techniques, and the lack thereof, is accentuated in the case of African languages, most of which are considered low-resource.
To help circumvent this issue, we explore techniques exploiting the qualities of morphologically rich languages (MRLs), while leveraging pretrained word vectors in well-resourced languages. In our exploration, we show that a meta-embedding approach combining both pretrained and morphologically-informed word embeddings performs best in the downstream task of Xhosa-English translation.

\end{abstract}
\section{Introduction \& Related Work}
Learning distributed representations of words \citep{mikolov2013distributed,pennington-etal-2014-glove} has found great success in many NLP tasks. 
However, the need for large amounts of data in low-resource African languages is a significant drawback of the current methods for learning word embeddings.
 \citet{shikali2019better} proposed learning Swahili word embeddings by taking morphemes into account, while \citet{grave2018learning} developed word embeddings for 157 languages, including multiple low-resource languages.

Learning of cross-lingual word embeddings \citep{DBLP:journals/corr/Ruder17} and projecting monolingual word embeddings to a single cross-lingual embedding space \citep{artetxe2018acl} have also been proposed to help learn embeddings for low-resource languages.  

In recent years, definitions and context words have been used to learn representations of nonce words (a word created for a single occasion to solve an immediate problem of communication), with \citet{doi:10.1111/cogs.12481} proposing learning nonce words by summing the representations of context words, obtaining representations highly correlated with human
judgements in terms of similarity. Additionally, \citet{TACL711} propose using word definitions to learn nonce word representations.

Our contributions are twofold: (1) Inspired by the work above, we explore techniques for combining pretrained high resource vectors and subword representations to produce better word embeddings for low-resource MRLs, and (2) we develop a new dataset, \textit{XhOxExamples}, made up of 5K Xhosa-English examples, that we collected from the \citet{oxford}.



\section{Approach}

Our approach assumes the existence of a vocabulary in our low-resource language $\sV = \{v_1,\dots,v_T\}$, the corresponding translations of the words in $\sV$ in a high-resource language, referred to as $\sD = \{d_1, \dots, d_T\}$, and a pretrained embedding matrix for the high resource language $E_{HR}$. $\sD$  
can either be comprised by individual words,
in the case the word in the low-resource language can be accurately mapped to a single word in the high-resource language (e.g. indoda $\rightarrow$ man)\footnote{All examples are from Xhosa to English}; 
or by a sequence of words in the case that the word in the low-resource language maps to more than one word in the high-resource language (e.g. bethuna $\rightarrow$ [listen, everyone]). Concretely, our objective is to use both $\sV$ and $\sD$ to produce vector representations for each word in $\sV$.

To leverage the high-resource language, we embed the atomic elements of $\sD$ in $E_{HR}$ and map the resulting vectors to the corresponding word in $\sV$. In the case of $d_i$ being a sequence, we take a similar approach to \citet{doi:10.1111/cogs.12481} and sum the normalized word vectors for each word in $d_i$ to produce a word representation for the word $v_i$. We refer to the resulting embedding matrix as $E_{V}$. Additionally, we pretrain another embedding matrix $E_{M}$ on a corpus in our low resource language using subword information to capture similarity correlated with the morphological nature of words \citep{DBLP:journals/corr/BojanowskiGJM16}.

We experiment with the following 4 methods to initialize the word embeddings for our downstream task\footnote{Note that the vocabulary for our downstream task might differ from $\sV$.}:
\begin{itemize}
    \item \textbf{\textit{Xh}Pre} - Initialization with $E_{V}$. Words not present in $E_V$ are initialized with $E_{M}$.
    \item \textbf{\textit{Xh}Sub} - Initialization with $E_{M}$ only.
    \item \textbf{VecMap} - We learn cross-lingual word embedding mappings by taking two sets of monolingual word embeddings, $E_{V}$ and $E_{M}$, and mapping them to a common space following \citet{artetxe2018acl}.
    \item \textbf{\textit{Xh}Meta} - We compute meta-embeddings for every word $w_i$ by taking the mean of $E_{V}(w_i)$ and $E_{M}(w_i)$, following \citet{coates-bollegala-2018-frustratingly}. Words not present in $E_V$ are associated with an UNK token and its corresponding vector.
\end{itemize}

\begin{table}
\begin{center}
\begin{tabular}{l|ll|llr} 
\toprule
& \multicolumn{2}{c|}{Bible}
& \multicolumn{2}{c}{XhOxExamples}\\
Aspect & Xhosa  & English &  Xhosa & English\\
\midrule
Sentence Length (Mean$\pm$Std)  & 18.44$\pm$7.74 & 29.43$\pm$12.61   &
5.16$\pm$1.89 &
8.42$\pm$3.04 \\
Total \# of Tokens & 573571 & 915531 & 25978 & 42400\\
\# of Sentences & \multicolumn{2}{c|}{31102}
& \multicolumn{2}{c}{5033}\\
Train/Dev/Test Ratio & \multicolumn{2}{c|}{70/20/10}
& \multicolumn{2}{c}{70/20/10}\\
\bottomrule
\end{tabular}
\end{center}
\caption{\label{tab:stats}Statistics regarding the Bible corpus and \textit{XhOxExamples} }
\end{table}
\begin{table}
\begin{center}
\begin{tabular}{l|ll|lr} 
\toprule
& \multicolumn{2}{c|}{Bible}
& \multicolumn{2}{c}{XhOxExamples}\\
Model & BLEU  & BEER & BLEU-4 & BEER\\
\midrule
Random Initializaton  &  21.79    & 21.84 & 16.08 & 25.30 \\
VecMap          &    22.46     &  22.03 & 16.38 & 25.42 \\
\textit{Xh}Sub &  24.65 & 22.79 & 17.37 & 26.04 \\
\textit{Xh}Pre       & 27.67     & 22.40 & 17.06 & 25.70\\
\midrule
\textit{Xh}Meta        & \textbf{29.09}  & \textbf{23.33}  & \textbf{17.77} & \textbf{26.44}   \\
\bottomrule
\end{tabular}
\end{center}
\caption{\label{tab:table}Results on the test set of the Xhosa-English Bible Corpus and \textit{XhOxExamples}}
\end{table}

\section{Experiments \& Results}

To test the performance of the 4 different embedding configurations, we both train and evaluate using machine translation (MT) as our downstream task on the English and Xhosa versions of the parallel Bible corpus \citep{christodouloupoulos2015massively} and make use of two MT evaluation metrics, BEER \citep{stanojevic-simaan-2014-beer} and sentence-level BLEU \citep{papineni2002bleu}. We also fine tune and evaluate the above pretrained models on the \textit{XhOxExamples} dataset to evaluate performance after fine-tuning as shown on the right side of Table \ref{tab:table}. Statistics for both datasets can be seen in Table \ref{tab:stats}.

All models are trained using OpenNMT \citep{klein-etal-2017-opennmt}, with a 2-layer 128-dimensional BiGRU for encoding, and 2-layer 128-dimensional GRU \citep{Cho_2014} using a beam search algorithm with a beam size of 5 for decoding. Embedding matrix $E_V$ is initialized with the 840B-token version\footnote{\url{https://nlp.stanford.edu/projects/glove/}} of GloVe \citep{pennington-etal-2014-glove}, while $E_M$ is initialized with a \texttt{FASTTEXT} model \citep{DBLP:journals/corr/BojanowskiGJM16}, trained on a combination of \textit{XhOxExamples} and the Xhosa Bible Corpus.

As is evident from the results shown in Table \ref{tab:table}, \textit{Xh}Meta outperforms all other models in both benchmarks. We believe this is a consequence of the simple approach of meta-embedding, resulting in vectors being both similar to both the subword representations captured by $E_M$ and the pretrained vectors, capturing a sense of the ``meaning" of the word, demonstrating the necessity of both aspects in this context.
\section{Conclusion}
In this paper, we explore different ways of combining pretrained high-resource word vectors and subword representations to produce more meaningful word embeddings for low-resource MRLs. We show that both types of representations are important in the context of the low-resourced MRL, Xhosa, and hope that this research is able to assist others when doing NLP for African languages.
\\~\\

\large{\textbf{Acknowledgements}}
\normalsize

We are grateful to Jorge Balazs for his fruitful
corrections and comments. We also gratefully acknowledge the support of NVIDIA Corporation with the donation of the Titan Xp GPU used for this research.
\bibliography{iclr2020_conference}
\bibliographystyle{iclr2020_conference}

\end{document}